# Parallel computations for Metropolis Markov chains with Picard maps

### v1

@@ A.8 Proof of Proposition 4
$=\frac{\sum_{i=0}^{\lfloor bd\rfloor}\mathbb{P}(f(X^{(j)}_{i},W_{i})\neq f(X_{i},W_{i}))}{\lfloor bd\rfloor r}$
@@
Appendix B Proofs for Metropolis within Gibbs
B.1 Proof of Theorem 3
Theorem 3 essentially shows that most of the results derived in Section 3.3 holds (with different constants) also for Metropolis within Gibbs (MwG) Markov chains, i.e. when replacing Assumption 3 with Assumption 4.
Lemmas 5-6 are the building blocks of all the results of Section 3.3. Thus, we start by providing the analogue of these Lemmas for MwG.
Lemma 14.
Under Assumptions 2 and 4, the statement of Lemma 5 holds.
To prove Lemma 14, we first need the following proposition.
Proposition 6.
Let $W^{\mathrm{MwG}}$ and $W^{\mathrm{RWM}}$ be the innovations of MwG and RWM satisfying respectively Assumption 4 and Assumption 2. Then, for every $x\in\mathcal{X}$
$\mathcal{L}(f_{i}(x,W^{\mathrm{MwG}}))$ $=\mathcal{L}(f(x,W^{\mathrm{RWM}}))$ for $\hbox{for }i=0,1,\dots,d-1,$
@@ B.1 Proof of Theorem 3
Proof.
By (8) and (11), Proposition 6 follows if $\mathcal{L}(o_{i}Z^{\mathrm{MwG}})=\mathcal{L}(Z^{\mathrm{RWM}})$ for all $i$, where $Z^{\mathrm{MwG}}=h/\sqrt{Ld}(2P-1)S$, $P\sim\mathrm{Ber}(1/2)$, $S\sim\chi(d)$, $Z^{\mathrm{RWM}}\sim\mathcal{N}(0,I_{d}h^{2}/(Ld))$, $(o_{0},o_{1},\dots,o_{d-1})\sim\mathrm{Haar}(d)$. By rotational invariance of $(o_{0},o_{1},\dots,o_{d-1})$ and $Z^{\mathrm{RWM}}$, it is enough to show that
@@
Proof of Lemma 14.
By Proposition 6, Lemma 5 follows also under Assumptions 2-4, with exactly the same constants and proof provided in Section A.2. ∎
Next, we provide the analogue of Lemma 6 for MwG.
Lemma 15.
Under Assumption 4 and for all $x,y\in\mathcal{X}^{K+1}$, with $x_{0}=y_{0}$, $w_{0}\in\mathcal{W}$ and $1<i\leq d$
@@
The proof of Lemma 15 relies on the following bounds.
Lemma 16.
@@
Proof.
By Laurent and Massart [2000, Lemma 1], for every $t\geq 2$,
@@
Then, Combining this result and Lemma 4, we have that
@@
with $\delta(d)=11\exp(-d/10)$. ∎
Proof of Lemma 15.
Under Assumptions 2 and 4,
@@
Since $\langle o_{i},o_{j}\rangle=0$ almost surely for all $i\neq j$ and by Proposition 6, we have
@@
with $\tilde{Z}_{1},\,\tilde{Z}_{2},\dots,\tilde{Z}_{i-1}\overset{\mathrm{i.i.d}}{\sim}\chi^{2}(d)$. Thus, by Lemma 16, we have
@@
with $\delta(d)=11\exp(-d/10)$. ∎
Proof of Theorem 3.
Under Assumptions 2 and 4, the statements of Lemma 8 and Theorem 1 hold with new constants
@@
The proofs follow the same steps of the proof of Lemma 8 and Theorem 1 in Section A.3.
Similarly, Under Assumptions 2 and 4, the statements of Corollary 1 and Theorem 2 hold with constants $c_{0},\delta(d)$ as in (23), $K=\epsilon\sqrt{\frac{d}{2c_{0}}}$ and $h\geq 5$. The proofs follow the same steps of the proof of Corollary 1 in Section A.5 and the proof of Theorem 2 in Section A.4.
Finally, Under Assumptions 2 and 4 the statement of Proposition 4 holds with $c_{0}$ as in (23) and $d\geq-10\log(\epsilon r/33)$. ∎
B.2 Proof of Proposition 3
Let $\pi(x)\propto\exp(-\|x\|^{2}/(2\sigma^{2}))$ and let $X^{(0)}_{i}=X_{0}$ for $i=0,1,\dots,K$. We prove here that $L^{(1)}=K$, for all $X_{0}=x_{0}\in\mathcal{X}$, $K\leq d$. Proposition 3 then follows by recursion.
@@ B.2 Proof of Proposition 3
$\langle\hat{z}_{\ell},\hat{z}_{i}\rangle$ $=z_{\ell}z_{i}\langle o_{\ell},o_{i}\rangle{=}0$ for all $\hbox{for all }i\neq\ell,$
@@
Appendix C Pseudocode
Algorithm 1 Online Picard algorithm
1:$X_{0}\in\mathcal{X}$, $N\in\mathbb{N}$, $K\in\mathbb{N}$.
2:Initialize $\bar{X}^{c}_{i}=\bar{X}_{i}\leftarrow X_{0}$, for $i=0,1,\dots,K$.
3:Set $L\leftarrow 0$
@@ Appendix C Pseudocode
5:while $L\geq N$ do
6:    $\bar{X}\leftarrow\Phi(\bar{X}^{c},\bar{W})$
7:    $G\leftarrow\sup\{i\leq K\colon\bar{X}_{\ell}=\bar{X}^{c}_{\ell}\text{ for }0\leq\ell\leq i\}$
8:    $\bar{X}^{c}_{0}\leftarrow\bar{X}_{G}$
9:    for $i=0,1,\dots,K-1$ do
10:       if $i\leq K-G-1$ then
11:          $\bar{X}^{c}_{i+1}\leftarrow\bar{X}_{G+i+1}$
12:          $\bar{W}_{i}\leftarrow\bar{W}_{G+i}$
13:       else
14:          $\bar{X}^{c}_{i+1}\leftarrow\bar{X}_{K}$
15:          $\bar{W}_{i}\sim\nu$
16:       end if
17:    end for
18:    $L\leftarrow L+G$
19:end while
20:return $\bar{X}_{G-(L-N)}$
Our implementation of the Online Picard algorithm is based on Proposition 1 and is provided in Algorithm 1. Its output is the random variable $X_{N}$, for any $N\geq 0$ given as input. The Approximate Online Picard algorithm is identical to Algorithm 1 except for line 5, which is replaced by
$G\leftarrow\sup\{1\leq i\leq K\colon\bar{\mathcal{A}}_{\ell}\leq r\text{ for all }\ell\leq i\},$

### v2

@@ B.2 Contraction of the Picard map: key lemmas
$\leq\frac{\sum_{i=0}^{\lfloor bd\rfloor}\mathbb{P}(f(X^{(j)}_{i},W_{i})\neq f(X_{i},W_{i}))}{\lfloor bd\rfloor r}$
@@
Appendix C Proofs for Metropolis within Gibbs
C.1 Proof of Theorem 4.1
Theorem 4.1 essentially shows that most of the results derived in Section 3.3 holds (with different constants) also for Metropolis within Gibbs (MwG) Markov chains, i.e. when replacing Assumption 3.2 with Assumption 4.2.
Lemmas B.9-B.11 are the building blocks of all the results of Section 3.3. Thus, we start by providing the analogue of these Lemmas for MwG.
Lemma C.1.
Under Assumptions 3.1 and 4.2, the statement of Lemma B.9 holds.
To prove Lemma C.1, we first need the following proposition.
Proposition C.2.
Let $W^{\mathrm{MwG}}$ and $W^{\mathrm{RWM}}$ be the innovations of MwG and RWM satisfying respectively Assumption 4.2 and Assumption 3.1. Then, for every $x\in\mathcal{X}$
$\mathcal{L}(f_{i}(x,W^{\mathrm{MwG}}))$ $=\mathcal{L}(f(x,W^{\mathrm{RWM}}))$ $\hbox{for }i=0,1,\dots,d-1,$
@@ Appendix C Proofs for Metropolis within Gibbs
Proof C.3.
By (5) and (8), Proposition C.2 follows if $\mathcal{L}(o_{i}Z^{\mathrm{MwG}})=\mathcal{L}(Z^{\mathrm{RWM}})$ for all $i$, where $Z^{\mathrm{MwG}}=h/\sqrt{Ld}(2P-1)S$, $P\sim\mathrm{Ber}(1/2)$, $S\sim\chi(d)$, $Z^{\mathrm{RWM}}\sim\mathcal{N}(0,I_{d}h^{2}/(Ld))$, $(o_{0},o_{1},\dots,o_{d-1})\sim\mathrm{Haar}(d)$. By rotational invariance of $(o_{0},o_{1},\dots,o_{d-1})$ and $Z^{\mathrm{RWM}}$, it is enough to show that
@@
Proof C.4 (of Lemma C.1).
By Proposition C.2, Lemma B.9 follows also under Assumptions 3.1-4.2, with exactly the same constants and proof provided in Section B.10.
Next, we provide the analogue of Lemma B.11 for MwG.
Lemma C.5.
Under Assumption 4.2 and for all $x,y\in\mathcal{X}^{K+1}$, with $x_{0}=y_{0}$, $w_{0}\in\mathcal{W}$ and $1<i\leq d$
@@
The proof of Lemma C.5 relies on the following bounds.
Lemma C.6.
@@
Proof C.7.
By Laurent & Massart (2000, Lemma 1), for every $t\geq 2$,
@@
Then, Combining this result and Lemma B.7, we have that
@@
with $\delta(d)=11\exp(-d/10)$.
Proof C.8 (of Lemma C.5).
Under Assumptions 3.1 and 4.2,
@@
Since $\langle o_{i},o_{j}\rangle=0$ almost surely for all $i\neq j$ and by Proposition C.2, we have
@@
with $\tilde{Z}_{1},\,\tilde{Z}_{2},\dots,\tilde{Z}_{i-1}\overset{\mathrm{i.i.d}}{\sim}\chi^{2}(d)$. Thus, by Lemma C.6, we have
@@
with $\delta(d)=11\exp(-d/10)$.
Proof C.9 (of Theorem 4.1).
Under Assumptions 3.1 and 4.2, the statements of Lemma B.15 and Theorem 3.1 hold with new constants
@@
The proofs follow the same steps of the proof of Lemma B.15 and Theorem 3.1 in Section B.3.
Similarly, Under Assumptions 3.1 and 4.2, the statements of Corollary 3.4 and Theorem 3.5 hold with constants $c_{0},\delta(d)$ as in (21), $K=\epsilon\sqrt{\frac{d}{2c_{0}}}$ and $h\geq 5$. The proofs follow the same steps of the proof of Corollary 3.4 in Section B.5 and the proof of Theorem 3.5 in Section B.4.
Finally, Under Assumptions 3.1 and 4.2 the statement of Proposition 5.1 holds with $c_{0}$ as in (21) and $d\geq-10\log(\epsilon r/33)$.
C.2 Proof of Proposition 4.2
Let $\pi(x)\propto\exp(-\|x\|^{2}/(2\sigma^{2}))$ and let $X^{(0)}_{i}=X_{0}$ for $i=0,1,\dots,K$. We prove here that $L^{(1)}=K$, for all $X_{0}=x_{0}\in\mathcal{X}$, $K\leq d$. Proposition 4.2 then follows by recursion.
@@
$\langle\hat{z}_{\ell},\hat{z}_{i}\rangle$ $=z_{\ell}z_{i}\langle o_{\ell},o_{i}\rangle{=}0$ $\hbox{for all }i\neq\ell,$
@@
Appendix D Efficient Implementation of the (Approximate) Online Picard Algorithm
Our efficient implementation is based on reformulating the Picard algorithm as a Markov chain on $K$-dimensional vectors.
Let $(X^{(j)},L^{(j)})_{j=0,1,\dots}$ be the sequence of random variables produced by the Online Picard recursion of (3)-(4). Define $\bar{X}^{(j)}=X^{(j)}_{L^{(j)}:U^{(j)}}$ and $\bar{W}^{(j)}=W_{L^{(j)}:U^{(j)}-1}$.
Proposition D.1.
Proof D.2.
@@ Appendix D Efficient Implementation of the (Approximate) Online Picard Algorithm
[…]
and (26)-(27) define the one-step Markov transition kernel of $(\bar{X}^{(j)},\bar{W}^{(j)})_{j=0,1,\dots}$.
Proposition D.1 sheds lights to the underlying (Markov) structure of Online Picard recursion and is used in Algorithm 3 for a memory-wise and computationally efficient implementation of the algorithm that iterates over a vector of length $K$. The Approximate Online Picard algorithm is identical to Algorithm 3 except for line 5, which is replaced by
$G=\sup\{1\leq i\leq K\colon\bar{\mathcal{A}}_{\ell}\leq r\text{ for all }\ell\leq i\},$
@@
Input: $N,K\in\mathbb{N}$, $X_{0}\in\mathcal{X}$.
Initialize $\bar{X}^{c}_{i}=\bar{X}_{i}=X_{0}$, for $i=0,1,\dots,K$;
Set $L=0$;
Set $\bar{W}_{0},\bar{W}_{1},\dots,\bar{W}_{K-1}\overset{\text{i.i.d}}{\sim}\nu$ ;
while $L<N$ do
 $\bar{X}=\Phi(\bar{X}^{c},\bar{W})$;
 $G=\sup\{i\leq K\colon\bar{X}_{\ell}=\bar{X}^{c}_{\ell}\text{ for }0\leq\ell\leq i\}$;
 $\bar{X}^{c}_{0}=\bar{X}_{G}$ ;
 for $i=0,1,\dots,K-1$ do
    if $i\leq K-G-1$ then
       $\bar{X}^{c}_{i+1}=\bar{X}_{G+i+1}$;
       $\bar{W}_{i}=\bar{W}_{G+i}$;
    else
       $\bar{X}^{c}_{i+1}=\bar{X}_{K}$;
       $\bar{W}_{i}\sim\nu$ ;
 $L=L+G$;
Output: $\bar{X}_{G-(L-N)}$.
Algorithm 3 Online Picard algorithm
Appendix E Numerical simulations for MwG
We run the same numerical experiments as in Section 6.4 for the (Approximate) Online Picard algorithms applied to Metropolis within Gibbs was applied. Figure 8 summaries the results. $d,N,K$ and $r$ are as in Section 6.4.
Figure 8: Performance of Online Picard algorithm ($\bar{X}$) and its approximate versions ($\bar{X}_{r},\,r=5\%,\dots,20\%$) applied to MwG, with target being the linear regression model E1 (top panels), logistic regression model E2 (middle panels) and Poisson regression model E3 (bottom panels).
Appendix F Definition of Effective Sample Size
The Effective Sample Size (ESS) of a test function $f\colon\mathcal{X}\to\mathbb{R}$ is defined as
$\text{ESS}(f)=\frac{1}{1+2\sum_{i=1}^{\infty}\rho_{i}(f)}$ (29)
where $\rho_{i}(f)$ is the lag-$i$ autocorrelation function of $\{f(X_{i})\}_{i=1,2,\dots}$. The Online Picard algorithm simulates a trajectory $(X_{0},\dots,X_{N})$ of the original Markov chain with $N/\hat{G}$ parallel iterations, resulting in an increase of ESS relative to the original sequential implementation of $\hat{G}$.
Appendix G Practical algorithmic aspects
G.1 Parallel Architecture of Picard algorithms
Here we discuss how the main processor interacts with the $K$ parallel processors at each iteration of the Picard map in (2). Our discussion focuses on RWM, but the same considerations apply to MwG.
Recall that, by Assumption 3.2, a RWM increment is given by
$f(x,z,u)=\mathbf{1}\left(\frac{\pi(x+z)}{\pi(x)}>u\right).$ (30)
Therefore, a naive implementation of the Picard map in (2) requires each processor to evaluate $\pi$ twice to compute the increment. This contrasts with an efficient sequential implementation of RWM, which requires a single evaluation of $\pi$ per iteration (Algorithm 4).
Input: Position $x\in\mathbb{R}^{d}$, evaluation of $\pi(x)$, Gaussian noise $Z$.
Compute $\pi(x+Z)$;
With probability $\min(1,\pi(x+Z)/\pi(x))$ set $x^{\prime}=x+Z$, otherwise set $x^{\prime}=x$;
Output: $x^{\prime}$, $\pi(x^{\prime})$.
Algorithm 4 One iteration of the sequential implementation of RWM
However, a careful look at (2) and (30) reveals a more efficient recursion for the Picard map, in which each parallel processor requires only a single evaluation of $\pi$ (Algorithm 5).
Input: $K\in\mathbb{N}$, $X^{(i-1)}_{0:K}$, $W_{0:K-1}$, $\pi(X^{(i-1)}_{0})$, $B^{(i-1)}_{j}:=\mathbf{1}(\pi(X^{(i-1)}_{j}+Z_{j})/\pi(X^{(i-1)}_{j})>U_{j}),\,j=0,1,\dots K-1$.
Compute in parallel: $\pi(X^{(i-1)}_{j}+Z_{j})$ for $j=0,1,\dots,K-1$;
Set $B^{(i)}_{0}=\mathbf{1}(\pi(X^{(i-1)}_{0}+Z_{0})/\pi(X^{(i-1)}_{0})>U_{0})$;
for $j=1,2,\dots,K-1$ do
 if $B^{{}^{(i-1)}}_{j-1}=0$ then
      Set $\pi(X_{j}^{(i-1)})=\pi(X_{j-1}^{(i-1)})$;
 else
      Set $\pi(X_{j}^{(i-1)})=\pi(X^{(i-1)}_{j}+Z_{j})$;
  Set $B^{(i)}_{j}=\mathbf{1}(\pi(X^{(i-1)}_{j}+Z_{j})/\pi(X^{(i-1)}_{j})>U_{0})$;
   Set $X^{(i)}_{j}=X^{(i)}_{j-1}+B^{(i)}_{j-1}$;
Set $X^{(i)}_{K}=X^{(i)}_{K-1}+B^{(i)}_{K-1}$;
Output: $\pi(X_{0}^{(i)})$, $X^{(i)}_{0:K}$ $B^{(i)}_{0:K-1}$.
Algorithm 5 Efficient implementation of the Picard map $\Phi$ in (2) for RWM
Note that, the parallel architecture used in Algorithm 5 is comparable to that used in parallel implementations of Multiple-try schemes (Glatt-Holtz et al., 2024): in both cases, the input is broadcast to $K$ processors, which evaluate $\pi$ at different locations in parallel; see Figure 9 for an illustration.
Figure 9: Diagram of one parallel iteration of the Picard map in Algorithm 5: the input is sent to $K$ parallel processors $P_{1},\dots,P_{K}$, each evaluating $\pi$ at a different location. A similar parallel architecture is used for parallel implementations of Multiple-try.
G.2 Parallelization overhead and effective speedup
To assess the effective speedup of Picard algorithms relative to their sequential counterparts, one must account for the wall-clock time of each algorithm. For a given algorithm, this mainly depends on (i) mixing of the Markov chain, as summarized by ESS in (29); (ii) the computational cost of each evaluation of $\pi$, denoted here by $c$; and, only for parallel algorithms, (iii) the parallelization overhead $\epsilon$ incurred by the parallel architecture used at each iteration. In this section we discuss how these factors affect the observed effective speedup.
We quantify efficiency via ESS per unit wall-clock time, i.e. $\mathrm{ESS}_{\mathrm{seq}}/T_{\mathrm{seq}}$ and $\mathrm{ESS}_{\mathrm{pic}}/T_{\mathrm{pic}}$ respectively for the sequential implementation and the online Picard algorithm. By the discussion in Section G.1, we let $T_{\mathrm{seq}}\approx N\times c$ and $T_{\mathrm{pic}}\approx M_{N}\times(c+\epsilon)$, where $M_{N}$ is the number of parallel Picard iterations needed to simulate $N$ increments of the Markov chain. The effective speedup of the Picard algorithm relative to the sequential algorithm is given by:
$\frac{\left(\mathrm{ESS}_{\mathrm{pic}}/T_{\mathrm{pic}}\right)}{\left(\mathrm{ESS}_{\mathrm{seq}}/T_{\mathrm{seq}}\right)}=\frac{T_{\mathrm{seq}}}{T_{\mathrm{pic}}}\approx\frac{N}{M_{N}}\left(1+\frac{\epsilon}{c}\right)^{-1}\approx\hat{G}\left(1+\frac{\epsilon}{c}\right)^{-1}.$ (31)
Here we used $\mathrm{ESS}_{\mathrm{pic}}=\mathrm{ESS}_{\mathrm{seq}}$, since both methods output the same trajectory. For instance, in our application in precision medicine (Section 6.8), $c\approx\epsilon$, i.e., parallelization overhead equals the cost of evaluating $\pi$, and the effective speedup is approximately half the speedup $\hat{G}$. Note that, by Theorem 3.5, we have that $\hat{G}=C\sqrt{d}$, when $K\geq\sqrt{d}$, for some $C>0$. Therefore, by (31), achieving an observed speedup of at least $S$ (with $S\leq C\sqrt{d}$) requires the parallelization latency $\epsilon$ to be smaller than a fraction $(C\sqrt{d}/S-1)$ of $c$. This condition can be checked for a given application and depends on the dimensionality of the problem, the target-evaluation cost $c$ and the parallelization overhead $\epsilon$, whose value is strongly influenced by the specific parallel architecture and hardware employed. For example, Glatt-Holtz et al. (2024) notes that, for Multiple-try, a GPU implementation with $K=10^{6}$ has roughly the same parallelization overhead of an equivalent implementation on a CPU with only $K=10^{2}$.